### v1

@@ 2.2 Small cases
defines an isomorphism ∎
The case is a little bit more complicated. We divide the analysis in a series of lemmas:
Lemma 2.4.
There is at least two classes in under relation
@@
Note that the matrices
If we assume that then by the second and third equations in 2.2.1, we have and but this contradicts the first equation in 2.2.1, therefore for all implying that is not invertible. ∎
Given a SLTM, we denote
Lemma 2.5.
Given a SLTM, such that and then we have
Proof.
1.
If then and the matrix
defines an isomorphism
2.
If then the matrix
defines an isomorphism
Lemma 2.6.
Given a SLTM, such that or then
Proof.
The following list shows all the different cases for satisfying the hypothesis of this lemma, and the corresponding matrix that defines an isomorphism
(Note that in the last case, we have since ). ∎
@@
The results provided in this section invite us to ask ourselves two natural question:
- Q1:
Is the number of classes in (under relation ) always finite?
We do not have an answer for those questions yet, but with the techniques developed we are able to find at least a lower bound for that number, see subsection 2.4 below.
@@ 2.3 The class of
In this section we describe a very explicit criterium (theorem 2.9) for determining when a given SLTM, is in the same class of the matrix This criterium is a generalization of lemma 2.6.
We first need the following definition:
Definition 2.8.
Given a SLTM, a scalar and indices we define the triangular subdeterminant
Note that if we take for we recover the number defined in subsection 2.2.
Let and the generators of and respectively.
Theorem 2.9.
Given a SLTM, then if and only if one of the following conditions holds
1.
for each pair of indices
2.
Given a SLTM, then for each pair of indices such that we have that for each
@@
for each and each Lets assume that there is a for some pair with Since the matrix is invertible, there is a such that therefore equation 2.3.1 implies that
@@
If the set is not empty, we say that it is a chain of leaders of If is a (not trivial) connected component of then we say that is a maximal chain of leaders. If are two different connected components of then we say that and are two independent chains of leaders of
@@
The proof of lemma 2.12 is direct, we left to the readers.
@@
1.
Let
(2.3.6) then In this case is itself a ramified chain of leaders of The associated graph looks like this:
Note that is composed by two simple chains: and
2.
Let
(2.3.7) Then In this case the associated graph is disconnected and it looks like this:
The sets are two independent simple chains of leaders such that
3.
Let
(2.3.8) In this case then is itself a simple chain of leaders. The associated graph in this case looks like this
4.
Consider the matrix
(2.3.9) In this case is itself a ramified chain, composed by two simple chains where
The associated graph in this case looks like this:
5.
@@
(2.3.10) Then the set is a ramified chain of leaders of The directed graph associated is given by:
Note that if we write and then are two (not independent) simple chain of leaders of such that
6.
Consider the matrix:
(2.3.11) Then the set is composed by two independent chains of leaders. The directed graph associated is given by:
Note that is divided in two independent (maximal) chains:
Also note that is a simple chain, while is a ramified chain, with where
@@
Let and suppose that In particular this implies that there is a such that By our hypothesis, for each and this implies that if and only if In particular, we have that This implies that in In fact, if we have that Otherwise, if then since there is a such that and again this implies that By a recursive argument, we can find a finite sequence: such that for each that is in other words
@@
Fix an If we assume that in then we have that:
The last equations imply that:
Therefore we have:
@@
Fix an . If there is a finite sequence: such that for each that is, then we have:
(2.3.12) (2.3.13) (2.3.14) From the equations in 2.3.12 we obtain respectively:
(2.3.15) Using recursively equations 2.3.13 and 2.3.14 we obtain respectively:
(2.3.18) and
(2.3.19) Finally we obtain:
as desired.
@@
Assuming that is itself a simple chain of leaders, say such that, for each we have for each Then
@@
First note that the hypothesis implies that the presentation of is given by the relations
also the hypothesis implies that the all the entries of are determined by the values of with Since is a chain we have that for each We also denote
We define for each minimal element (in particular for ), the element Being minimal means that there is no such that in particular by lemma 2.12, we know that
@@
One can check recursively (similarly to the proof of lemma 2.14) that:
that is
and so on.
@@
and
Now note that condition for each implies that
@@
for any
In fact, it is clear that
Assuming that equation 2.3.25 is true, for any then we have:
therefore
and equation 2.3.25 is true for any for any
Example 2.16.
Let be the matrix of equation 2.3.8, then then the presentation of is given by the equations
@@
For not minimal, we define recursively:
By property 2.3.4 of lemma 2.12, is well defined by equation 2.3.26.
By item 4 of lemma 2.12, we know that the set is a simple chain of leaders, therefore the fact that follows analogously to the proof of this in lemma 2.15. Note that each independent chain can be treated independently. Therefore we have that the assignment defines an homomorphism
It is clear that the associated matrix is invertible, therefore is an isomorphism. ∎
@@
Example 2.20.
Let be the matrix of equation 2.3.9. As we mentioned in example 2.13 is a ramified chain, composed by two (not independent) simple chains, and One can check that satisfies all the conditions of proposition 2.17. The isomorphism described in the proof of proposition 2.17, is given by:
The associated matrix is
Example 2.21.
Let the matrix of equation 2.3.10. As we have mentioned in example 2.13, is a ramified chain composed by the two (maximal) simple chains: and Note that in this case the minimal vertices are all of them but are isolated vertices of Also note that satisfies the hypothesis of proposition 2.17. Therefore
Following the proof of proposition 2.17, we have that:
Using the recursive definition provided in equation 2.3.26 in the proof of proposition 2.17, we have for :
Analogously for we obtain:
@@ 2.4 A lower bound for the number of classes
In this subsection we show the following:
@@
To do so, we shall show that is a set of representatives of pairwise different classes under relation
Lemma 2.24.
Considering the matrix we have that if and only if
@@
Is not difficult to see that the matrix satisfies the conditions of theorem 2.9 if and only if ∎
Proposition 2.25.
The matrices and are related under if and only if
Proof.
By lemma 2.24, we know that whenever Therefore we only have to prove that the matrices and for are not related by
If we assume that there is an isomorphism then by theorem 2.2, we have
@@
If we assume that there is a such that then equation 2.4.2 implies that whenever and On the other hand, if we assume that for all then necessarily since But if we replace in equation 2.4.1, we obtain and therefore for all but this is a contradiction with the definition of the matrix The last analysis implies that for each there is a such that and for each with
Now we replace in equation 2.4.1, and such that then we obtain Since by definition of the matrix then we conclude that That is valid for each Therefore, since we conclude that By the last analysis we also conclude that for each pair we have In particular, we conclude that there are exactly rows such that
If we replace in equation 2.4.1 we obtain
and by definition of the matrices and the indices we obtain
Note that by the symmetry of equation 2.4.3 we can omit the hypothesis that
Going back again to equation 2.4.1 but this time taking we obtain
By definition of the matrix and by the facts that for each and we obtain more precisely that
If we take for some we obtain
Let us fix such that (we know there is at least two rows satisfying that condition. Then we have:
@@
The last two cases implies that implies that and then there are at most rows such that but this is a contradiction, since we know there are exactly rows satisfying that condition. By the last analysis we conclude that and are not isomorphic, and then ∎
2.5 Elementary triangular operations
In this section we define a series of matrix operations called Elementary triangular operations (ETO), that preserves the class under relation At the end of the subsection we prove for any pair that for having is necessary and sufficient that there exists of a (finite) sequence of ETO that transform into
Definition 2.26.
Let be a SLTM, fix an and an scalar We define the matrix as the SLTM, where
Example 2.27.
Consider the matrix
then
Lemma 2.28.
Let ba a SLTM, fix and an scalar Then
Proof.
Let Let and the generators of and respectively. We claim that the assignment
defines an isomorphism To see that recall that by equation 2.1.3, we only have to check that and the associated matrix is invertible. Now note that
1.
If then clearly
2.
If then we have
and
3.
if then we have
and
By the last analysis, we conclude that the assignment defines an homomorphism Is not difficult to see that the associated matrix is invertible, in fact Therefore we conclude that defines an isomorphism ∎
Definition 2.29.
1.
The entry for each
2.
The entry for each
then we define the matrix as the SLTM, where
Example 2.30.
Consider the matrix
and take and Is not difficult to see that satisfies the conditions of definition 2.29. Then the matrix is well defined. Moreover we have
Lemma 2.31.
Let be a SLTM, let If satisfies the conditions of definition 2.29, then
Proof.
Let Let and the generators of and respectively. We claim that the assignment
defines an isomorphism To see that, note that by our hypothesis we have:
1.
For we have
and
2.
For we have
and
3.
For it is immediate that
4.
For we have
and
5.
For we have
and
The last analysis implies that the assignment defines a homomorphism Is not difficult to see that the associated matrix is invertible, in fact Therefore defines an isomorphism ∎
Recall that for a given SLTM, we denote
Definition 2.32.
1.
2.
and for each
Then we define the matrix as the SLTM, given by
Example 2.33.
Consider the matrix
Example 2.34.
Therefore we have
Lemma 2.35.
Proof.
Let Let and the generators of and respectively. We divide the analysis in two cases:
1.
If then we have that the matrix is given by:
We claim that the assignment:
defines an isomorphism In fact, if then it is clear that
We also can check for that
and
Finally for we have
and
Therefore, the assignment defines a morphism Is not difficult to see that the matrix is invertible, moreover then is an isomorphism.
2.
If and for each then the matrix as the SLTM, is given by
We claim that the assignment:
defines an isomorphism In fact, if then it is clear that
We also can check for that
On the other hand
Note that by our hypothesis, we have then
then as desired.
Finally for we have
and
Therefore, the assignment defines a morphism Is not difficult to see that the matrix is invertible, moreover then is an isomorphism.
Definition 2.36.
We call Elementary triangular operations (ETO for short) any of the following matrix assignments (whenever they are well defined):
Given two SLTM, and we say that they are equivalent by ETO if there is a finite sequence of ETO that transform into We denote by whenever and are equivalent by ETO.
Example 2.37.
Consider the matrices
then In fact one can check that
Definition 2.38.
In the following we shall denote respectively by the matrices given by:
Note that the matrix described in the proof of lemma 2.28 is of the form the matrix described in the proof of lemma 2.31 is of the form and the matrix described in the proof of lemma 2.35 is of the form
Theorem 2.39.
Let be two SLTM, then if and only if
Proof.
By lemmas 2.28, 2.31 and 2.35 we have that if then Moreover lemmas 2.28, 2.31 and 2.35 implies that for two SLTM, we have that if and only if there is an invertible matrix that can be written as a (finite) product of matrices of the form and
If then there exist an invertible matrix satisfying equation 2.1.6. Since is invertible, it can be factored as a product of elementary matrices (in the usual sense), Now the result follows from the fact that the matrices and are elementary matrices and any other elementary matrix can be written as a product of matrices of the form and ∎
3 Applications To Soergel Calculus
Fix a Coxeter system of finite rank (with for certain set of indices ), and a field where the number has sense in for each pair Following [6], the geometric realization of over a is a triple composed by a vector space subsets and where:
and the equations:
define representations of the group on and respectively. For any and any let
We are avoiding a technical definition on the diagrammatic Soergel category since it is well known. We invite the readers to look for more information in the original work of Elias and Williamson [3]. In this section we shall assume that the ground field has characteristic different from and satisfies all the technicalities described in [3]. Let the free monoid generated by the symbols In [12], for each element we define the SLTM:
and the corresponding nil graded algebra We have proved in [12], that the algebra corresponds to the Gelfand-Tsetlin subalgebra of the endomorphism algebra in the diagrammatic Soergel Category of Elias and Williamson (see [3]), relative to the Libedynski’s light leaves basis (see [8]) and the set of Jucys-Murphy elements defined by Ryom-Hansen in [19].
In [12] we show that for the case of a Coxeter system of type the algebras are invariant under commuting moves, and we proved that by defining explicit isomorphisms. Now in theorem 3.2 we generalize this fact for any other type. In this case we use the techniques developed in this article, to obtain a cleaner proof.
Let us recall first, certain technical definitions coming from [12]:
Definition 3.1.
1.
Let we define the column vector associated to as the matrix given by:
2.
Given two elements of say and we define the matrix by
where
In [12], we show that for any composed expression we have that:
Moreover in [12], we proved that for we have the following (kind of) associative property:
where
Now we are ready to prove the following:
Theorem 3.2.
For any Coxeter system the algebra is invariant under commuting moves.
Proof.
Let such that that its, the elements and of the group commute. Let and then are elements of that differ by a commuting move.
1.
Let and Then and Therefore:
(3.0.3) where
2.
Note that by definition and Therefore we have:
By the same reason, we have that two first rows of are given by:
analogously the two first rows of are given by:
3.
Note that for every we have that Let us denote
then
4.
Following with the notation of the s, we have that the rows of the matrices from the third to the last one are given respectively by:
and
But since and commute, we clearly have that
- 5.
Theorem 3.2 naturally opens the following question:
- Q3:
Is the algebra invariant under other braid moves?
Fortunately we have an answer for Q3, although it is not a positive one:
Proposition 3.3.
The algebra is not invariant under braid moves.
Proof.
It is enough to provide an example where two equivalent by braid moves expressions and generate two not isomorphic algebras and
Let be the Coxeter system of type Let and then:
1.
It is clear that and are equivalent by braid moves.
2.
The corresponding matrices and are given by:
3.
Considering the matrices defined by:
one can check that they are well defined (see subsection for details) and by definition More specifically we have:
4.
Let be the entries of let be the entries of the matrix By theorem 2.39, if the algebras and were isomorphic, then necessarily there will be an isomorphism Assuming that, then the associated matrix would satisfy equation 2.1.6 of theorem 2.2. In particular we would have for and any pair that:
(3.0.4) -
5.
Equation 3.0.4 in particular implies that:
- 6.
7.
Let and be the sets of generators of the algebras and respectively. From the last two items, we conclude that if there exists an isomorphism then necessarily for In particular this implies that the matrices and are in the same class under relation but this contradicts theorem 2.9. In fact
The analysis above implies that and are not isomorphic. ∎

### v2

@@ 3.1 Small cases
In the smallest case, that is, we can easily count exactly the number of classes:
@@
defines an isomorphism β
Corollary 3.2.
If are any two matrices in If then
@@
The case is a little bit more complicated, but again, we can count exactly the number of classes. For the rest of this subsection for a given SLTM, we denote (for notations, see subsection 1.2).
@@
We divide the proof in a series of steps:
- Step 1
Note that the matrices
(3.1.3) If we assume that then by the second and third equations in 3.1.3, we have and but this contradicts the first equation in 3.1.3, therefore for all implying that is not invertible.
Note that Step 1, implies that there are at least two classes in in under relation
- Step 2
- Step 3
Given a SLTM, such that or then In fact, the following list shows all the different cases for satisfying the hypothesis of this lemma, and the corresponding matrix that defines an isomorphism
One can check in each case that is an invertible matrix and that its entries satisfy equation 2.2.6 of Theorem 2.6, (taking and ). Note that in the last case, we have since
- Step 4
Let be a SLTM. By Step 2 and Step 3, we have that is in the class of or in the class of Finally Step 1 implies that there are exactly two classes in
β
Corollary 3.4.
If are any two matrices in If then
Proof.
Let be a SLTM. By the proof of Proposition 3.3, we have that:
β’
If then one can check that the matrix defined in Step 2 can be factored as follows:
that is
(3.1.4) Now this factorization yields the following well-defined process of transformations:
(3.1.5) Is not difficult to check that and therefore one conclude that
β’
If then one can check that in all the cases described in Step 3, one have In fact all the corresponding matrices can be factored in a product of matrices of the form and the factorization yields a well-defined process of transformations, via operations, therefore implies that We left details to the readers.
Let such that As a consequence of the proof of Proposition 3.3, this implies that one and only one of the following is true or Finally the above analysis implies in both cases that whenever β
@@ 3.2 The class of
We would like to understand, how the information coming from the codifying matrices, can be used to describe completely an specific isomorphism class in In this direction, in this section we describe a very explicit criterion (Theorem 3.12) for determining when a given SLTM, say is in the same class of the matrix This criterion is a generalization of Step 3 in the proof of Proposition 3.3.
Along this subsection we denote and the generators of and respectively. Also recall from subsection 1.2, the notation (or simply ) given by
@@
for each and each Let us assume that there is a for some pair with Since the matrix is invertible, there is a such that therefore equation 3.2.1 implies that
@@
If the set is not empty, we say that it is a chain of leaders of If is a (non-trivial) connected component of then we say that is a maximal chain of leaders. If are two different connected components of then we say that and are two independent chains of leaders of
@@
The proof of Lemma 3.7 is direct and it is left to the readers.
@@
Then the set of leaders is The directed graph associated is given by:
Note that is divided in two independent maximal chains of leaders, where
Also note that is a simple chain, while is a ramified chain, such that where are the simple chains of leaders given by:
Finally note that the minimal vertices of are , this is because the st, the th and the th rows of are zero row. In particular the vertex is an isolated vertex.
@@
Let and such that By our hypothesis, for each and this implies that if and only if In particular, we have that since
Fix an We have two cases:
@@
If we assume that (that is ), then by our hypothesis, we have that:
Then we have:
therefore
@@
If then since then necessarily there is a such that and again this implies that By a recursive argument, we can find a finite sequence: such that for each that is (in other words ).
In this case, by our hypothesis, we have:
(3.2.7) (3.2.8) (3.2.9) From the equations in 3.2.7 we obtain respectively:
(3.2.10) Using recursively equations 3.2.8 and 3.2.9 we obtain respectively:
β
@@
Assuming that is itself a simple chain of leaders. If for each and each we have then
@@
Recall that and denote the generators of and respectively. Let then the hypothesis implies that the presentation of is given by the relations
Since is itself a simple chain, we have that for each We also denote We define for each minimal element (in particular for ), the element Being minimal means that there is no such that in particular by Lemma 3.7, we know that
@@
Recall that we want to prove that then we need to examine the two terms on the right of equation 3.2.17 to see them clearer:
On one hand, one can check recursively (similarly to the proof of Lemma 3.9), that:
@@
On the other hand, note that condition for each implies that
@@
for any In fact, assuming that equation 3.2.20 is true, for any and noticing that that we obtain:
Therefore equation 3.2.20 is true for any for any
@@
By item 4 of Lemma 3.7, we know that the set is a simple chain of leaders, therefore the fact that follows analogously to the proof of this in Lemma 3.10. Note that each independent chain can be treated independently. Therefore we have that the assignment defines an homomorphism It is clear that the associated matrix is invertible, therefore is an isomorphism. β
Theorem 3.12.
Given a SLTM, then if and only if one of the following conditions holds
- 1
for each pair of indices
or
- 2
For each with we have that for each
Proof.
@@
Let be the matrix of equation 3.2.6. Note that satisfies the hypothesis of Proposition 3.11. As we have mentioned in Example 3.8, is divided in two independent chains, a ramified one with and a simple one
In this case the minimal vertices of are where only is isolated. Following the proof of Proposition 3.11, we can define the isomorphism by taking:
For minimal vertices:
For vertices associated to chain
For vertices associated to chain
And finally, for vertices associated to chain
The following Proposition provides another evidence of the veracity of Conjecture 2.21.
Proposition 3.14.
Let be a SLTM. If then
Proof.
Assuming that but We shall prove that by induction on the number of non-zero rows of
β’
If let be the leader of the only non-zero row of Since Theorem 3.12 implies that either or for each Both cases implies that is the only nonzero entry of and that the ETO: is well-defined over Is not difficult to see that and therefore
β’
If let the minimal value such that the th row of is different from zero. Let be the leader of the th row. Since Theorem 3.12 implies that either or for each Both cases implies that is the only nonzero entry of and that the ETO: is well-defined over Let then most of the entries of are equal to the corresponding entry of with the exception of those in positions of the form with such that For those cases we have the relation:
(3.2.22) (see equation 2.3.11). Therefore in the aforementioned positions, we have that
If is such that and is the leader of the th row of then since In particular, we have that and by definition of we have that By equation 3.2.22, we conclude that and then For any other triple with the entries of are the same as the corresponding entries of then it is clear that for each and each Theorem 3.12, implies that
By construction and by equation 3.2.22, is a matrix with non-zero rows. By induction we conclude that and by transitivity of relation we conclude as desired.
β
@@
Let be the matrix defined as the following product of elementary matrices:
where
One can check that the matrix defines an isomorphism and the factorization given in equation 3.2.23 yields a well-defined process of transformation via operations, that transform into This process can be deduced from the proof of Proposition 3.14.
Note that for the isomorphism defined in example 3.13, the associated matrix can also be factored as an adequate product of elementary matrices, that yields a well-defined process of transformation via operations, that transform into We left details to the readers.
@@ 3.3 A lower bound for the number of classes
We would like to know, how many classes can be found in , for each and how this number depends (or not) on the ground field Let us denote the number of classes in under As we observed in Subsection 3.1, the number for is finite and do not depend on the ground field. Although, at this point we do not know if this number is always finite or if it is independent or not on the ground field, in this subsection we obtain a lower bound for that is still independent on the ground field:
@@
We divide the proof in several steps, where basically we shall show that the matrices are representatives of pairwise different classes.
- Step 1
First note that if and only if In fact, note that the matrix satisfies the first condition of Theorem 3.12, then On the other hand, if then by definition of matrix (see Subsection 1.2), the entry in position is not zero and therefore the matrix do not satisfy the necessary condition of Theorem 3.12, and then as desired.
- Step 2
Let and with If we assume that there is an isomorphism then by Theorem 2.6, we have
(3.3.1) - Step 2.1
@@
(3.3.2) By our assumption, necessarily for each there is a such that (Since the associated matrix has to be invertible). If we assume that then equation 3.3.2 implies that whenever and On the other hand, if we assume that and we replace in equation 3.3.1, we obtain and therefore for all but this is a contradiction with the definition of the matrix The last analysis implies that for each there is a such that and for each with
Now we replace in equation 3.3.1, and such that then we obtain Since by definition of the matrix then we conclude that That is valid for each Therefore, since we conclude that
By the last analysis we also conclude that for each pair we have In particular, we conclude that there are exactly indices such that
- Step 2.2
Let us fix now a pair such that If we replace in equation 3.3.1, we obtain
(3.3.3) and by definition of the matrices and the indices we have that whenever and whenever , therefore equation 3.3.3 reduces to
(3.3.4) Note that by the symmetry of equation 3.3.4, we can omit the hypothesis that
- Step 2.3
If we replace in equation 3.3.1, we obtain
(3.3.5) By definition of the matrix (in particular using that if and if ), and by the facts that for each and we obtain, after simplification by that
(3.3.6) - Step 2.4
Let us fix If we replace in equation 3.3.6, then we have that and then we obtain
(3.3.7) - Step 2.5
Let us fix with such that (we know there is at least two of them satisfying that condition), in particular Therefore we have the following two cases:
@@
The last two cases imply that whenever we necessarily have Then there are at most indices such that but this is a contradiction, since we know from Step 2.1, that there are exactly of them satisfying that condition (recall that ).
By the last analysis we conclude that and are not isomorphic, and then whenever
- Step 3
By Step 1 and Step 2, we conclude that the matrices are representatives of pairwise different classes under relation Therefore there are at least different classes in .
β
- Q1:
Is the number always finite?
- Q2:
Is the number independent of the ground field?
We do not have an answer for this questions yet, but motivated by our intuition, we define the following:
Conjecture 3.17.
The number for depends on the ground field. Moreover the number is finite if and only if the ground field is finite.
We expect to be able to prove or disprove Conjecture 3.17 in future works.